OpenStudy (doublea124):

y^1/2 − 9y^1/4 + 20 = 0

OpenStudy (agent0smith):

Looks like a quadratic...\[\Large y^{1/2} − 9y^{1/4} + 20 = 0 \] \[\Large (y^{1/4})^2 − 9(y^{1/4}) + 20 = 0\]If you can't see it, do a substitution, replace y^1/4 with u: \[\Large u^2 − 9u + 20 = 0\]Now factor it and solve (don't forget to replace u with y^1/4)

OpenStudy (doublea124):

Thank you. This'll teach me never to skip class again, even if it's for another. The fractional exponents are what's throwing me off.